# Nonlinear Impulse Pattern Formulation dynamical social and political prediction algorithm for city planning and public participation

### v1

@@ III.1 Single stakeholder without self-adaptation with external disturbance
Five cases are shown in Fig. 3. The top blue curve shows $g^{1}$ over 50 iterations with $I_{t}=0$ without impact. After an initial transient, $g^{1}$ becomes stable and converges.
@@
The yellow curve shows the reaction to a positive $I_{20}=1$, so an impact of one at a single iteration time 20. The system is disturbed and returns to its converging state. This is expected as such a distortion corresponds to a new initial $g_{0}$ after which convergence is ensured due to the choice of $\alpha=0.56$ for which such a system is stable.
The green curve shows a stronger impact of $I_{20}=5$. Although the system becomes stable again, it takes more time to recover.
@@
All these behaviors represent reasonable reactions of a social system to an impact; it might converge to a stable value, even with a negative impact, or it might break down and leave the stage.
@@ III.2 Single stakeholder with self-adaptation with external impact
To estimate the effect of self-adaptation, the model again with s=1 and a strong impact at iteration time 20 of $I_{20}=10$ is used, but now including self-adaptation, so using varying term II in version I. Fig. 4 compares $g^{s}_{t}$ for both cases, starting at $\alpha=0.56$. In the above case, $\alpha$ = const, while now $\alpha$ is adaptive, therefore time-varying. Also now, $C^{s}=-1$ is used.
The green curve in Fig. 4 is $g^{s}_{t}$ of varying term I version II already shown in Fig. 3. The blue curve is the corresponding curve of varying term II version I with adaptive $\alpha$, shown in yellow. Both, $g^{s}_{t}$ and $\alpha_{t}$ are bifurcating but are stable in this respect. The impact at iteration time point 20 disturbs both curves considerably. Still, $g^{s}_{t}$ of version I returns to its bifurcating state much faster than $g^{s}_{t}$ of version II.
@@
Figure 4: Adaptive version I of varying term II compared to non-adaptive version II concerning reaction of external input at iteration time point 20. Green: Version II time series from Fig. 3 with constant $\alpha=0.56$. Blue: Version I time series with adaptive $\alpha$. Compared to Version II, the Version I time series is bifurcating due to adaptivity; still, it is returning to this state faster compared to Version II with static $\alpha$. Yellow: Version I adaptive $\alpha$ corresponding to Version I system parameter $g^{s}_{t}$. After the impact on time point 20 $\alpha$ is way above the diverging limit of $\alpha=0.37$, the system remains stable due to adaptation.
@@
All calculations were performed using an initial system value $g^{s}_{0}=0.1$. This ensures $\Im{g_{1}}=0$ for $\alpha_{0}>0.09048$ within the first iterative step, a value of $\alpha$ much lower than expected in the system at all. Using $g_{0}=1$ would lead to $\Im{g_{1}}\neq 0$ for $alpha_{0}\geq e$. Therefore, any adaptation algorithm would no longer be able to converge $\alpha_{0}<0.3679$, and extended stability by including any adaptation algorithm could not be tested. Using any $g_{0}$ is artificial in any case, as the initial time point of a social or political situation is arbitrarily set by an investigator while the stakeholders taking part in the social action most often existed before the chosen model starting time point.
@@
Fig shows the three cases for I = 0, so without adaptation, I = 1, and I = 10. The adaptation constant was varied from -3 $\leq C^{s}\leq 7$ and the initial $\alpha_{0}$ like 0.1961 $\leq\alpha_{0}\leq$ 10.
In the top plot of Fig. 5, colors denote converging cases of $C^{s}$ vs. $\alpha_{0}$ combinations where the color corresponds to the mean converged value. A tail appears starting from $\alpha_{0}$ = 10 and 0 $\leq C^{s}\sim$2 which is getting narrower towards $\alpha_{0}\sim$ and $C^{s}\sim$ 0.2. Therefore, choosing a suitable $C^{s}$ allows for much higher $\alpha_{0}$ to converge compared to the non-adaptive Version II with a minimum of $\alpha\sim 0.37$. Indeed, with decreasing $\alpha_{0}$, the range of $C^{s}$ for converges is getting smaller, as expected. Still, it is interesting that the convergence value is smaller towards higher $\alpha$ displayed by the lighter color.
The convergence area converges for all parameter values within the area and is not fragmented in the sense of non-converging islands. So, convergence is ensured within the displayed regions.
@@
An unsystematic convergence area also exists for $C^{s}<-2$, which is quite unstable. A stakeholder in this region might consider it quite random to maintain himself stable.
@@
III.3 Multiple stakeholders with direct impact and no external disturbance
Adding the interaction between multiple stakeholders leads to a temporal description of the system parameter $g^{s}$ for all stakeholders in the process. This is implemented by the interaction parameter $\beta^{s,i}_{t}$ of stakeholders i impacting on stakeholder s. First, direct impact is discussed, which means including the Variation I of Varying term III as shown in Tab. 1.
@@ III.3 Multiple stakeholders with direct impact and no external disturbance
In this simulation, we assume all stakeholders are already present in the system before the simulation starts at $t=0$. Still, couplings $\beta^{s,i}$ are often not known or hard to determine, which leads to the simplification of uncoupled stakeholders before $t=0$. Thus Eq. 1 and 5 are valid $\forall\ t>=0$. $\forall\ t<0$ the system can be described by the IPF with $\beta^{s,i}=0\forall\ s,i$. As we assume the system to be running for a sufficiently long time, all stakeholders have already converged to their fixed point. When assuming that no $\beta$ are involved before $t=0$ this fix point is $g^{s}_{f}=\alpha$ according to Eq. 1 and thus $g^{s}_{t}=\alpha^{s}\ \forall t<0$. Then, all necessary initial values of the simulation are set. Further, the time delay $d^{s,i}=i$ directly depends on the couped stakeholder.
In the following, IPFs with $N=6$ stakeholders are used throughout, as taken from the Frappant example. Each IPF is iterated for 500 steps using Equations 1 and 5. $d^{s,i}=1$ is used throughout. By randomly varying $\alpha^{s}$, IPFs were calculated until 1000 non-diverging time series were obtained. This amount was reached after 3523 attempts. Thereby, $\alpha^{s}$ was chosen within a range of $\alpha^{s}\in[0.2,1]$. Also, the coupling constant $W^{s}_{i}=W$ was fixed for all stakeholders but varied for each IPF again randomly in the range of $W\in[0.01,1/6]$.
@@
Figure 6: Trained SOM of 1000 IPF cases of six stakeholders using direct interaction with Variation I of Varying terms II. The background color is the u-matrix displaying similarity (dark colors) of differences (light colors) of neighbouring neurons. Therefore, the yellow ridges are boundaries between clusters present in darker regions. Each dot is an IPF $g^{s}$ iteration’, which is placed according to its best fit to the feature vectors of the trained SOM. The colors of the dots refer to the coupling strength $W$; darker colors are stronger, and lighter colors are weaker coupling. For each of the clusters or regions I to VIII, an assigned time series $g^{s}_{t}$ is picked and plotted aside the SOM as a typical example of stakeholders $g^{s}$ development of the last 20 values of the 500 iteration steps for the respective regions I to VIII.
Fig. 6 shows the trained SOM for Variation I of Varying term III. The background color is the so-called u-matrix, displaying similarities and differences of neighboring neurons. Lighter yellow colors have strong differences, and darker blue colors have small differences. Therefore, the yellow ridges are bounders between blue regions of very similar neurons. The dots represent the 1000 IPF cases the map was trained with, and they are positioned after training at the best-fit neurons for each IPF. The colors of the dots show varying coupling strengths $W$, where lighter colors are weaker and darker colors have stronger coupling.
@@
Figure 7: Example IPF time series of region I of Fig. 6 of $g^{s}$ converging for all s=6 stakeholders, calculated from Eq. 5. This example was initial $\alpha^{s}\approx\{0.98,0.71,0.79,0.75,0.55,0.77\}$ and $W\approx 0.011$. Therefore, the largest region I of all VIII regions is that of convergence of all stakeholders.
@@
A more systematic view of regions II - VII obtained when looking at Fig. 10. Here, for each node of the SOM, the last ten values of the time series $g^{s}_{t}$ are inspected in terms of the number of different values. Values are assumed to be different when they differ more than 5 % from each other. One can clearly distinguish the stable region I and the chaotic region VIII. Furthermore, II - VII, while looking rather similar in Figure 6, are found to differ in terms of the order of bifurcation.
Figure 10: Amount of different values of neurons of trained SOM of Fig. 6, when inspecting the last ten values of each stakeholder time series $g^{s}$. Region I clearly has the lowest values as it is converging throughout. Also, the chaotic region VIII has the highest values. Still, region II - VII can now be differentiated with respect to the number of different values and, therefore, the order of bifurcation.
Using only version I of varying terms III of direct interaction between stakeholders shows all kinds of possible scenarios for city planning. These include
  • Convergence: All stakeholders arrive at their desired value.
@@
So, the inverse problem of finding parameters and therefore, political actions to arrive at a desired output for all stakeholders is solved using the proposed method. Still, a second interaction process will appear between stakeholders, that of mediated interactions discussed below.
III.4 Multiple stakeholders with mediating impact and no external disturbance
Along the method of direct interaction discussed above, for mediated interaction again IPFs for $N=6$ stakeholders with 500 iteration steps each were calculated, using Equations 1 and now Version II of Varying terms III of Eq. 6.
@@ III.4 Multiple stakeholders with mediating impact and no external disturbance
  • I1: Same as the previous modeling example of direct impact, it is assumed that no coupling is present $\forall t<0$ and that the system has already settled itself, and thus $g^{s}_{t}=\alpha^{s}\ \forall t<0$. At $t=0$, suddenly, all stakeholders are coupled to each other.
  • I2: Similar to I1, but as the system is supposed to be uncoupled $\forall t<0$, it is not realistic that all stakeholders are suddenly coupled to each other. The IPF models the influence of coupled stakeholders by the damped influence of a preceding system states $g^{s}_{t-d^{s,i}}$. As coupling is only present after $t=0$, previous system states can not act on future ones. Thus, the different $\beta^{s,i}_{t}$ are only taken into account if $t\geq d^{s,i}$.
  • I3: It is assumed that the system is not settled at the start. Thus $g^{s}_{t}$ my have an arbitrary value $\forall t<0$. Here, for simplicity, we chose $g^{s}=1$. Like in I1 it is assumed that all stakeholders are suddenly coupled to each other at $t=0$.
@@
The resulting time series are shown in Fig. 13 to 16 as examples for each of the four initial conditions. They are very much different from those of the direct interactions discussed above. Bifurcations and chaos can rarely present, and, due to the mutual adaption of $\beta^{s,i}_{t}$, all coupled IPFs converge to the same fixed point. This fixed point might be reached in an exponential development, like shown in Fig.15, it might show vibrations before convergence, like in Fig. 13 - 14, or it might show very slow convergence, so basically each stakeholder keeps its own state.
An example is a legally binding land-use plan process (B-Plan Verfahren) in which all stakeholders negotiate over time on a planning of a area. The result is planning certainty. Due to negotiations, a mediating and self-reference process is leading to perfect convergence of all stakeholders needs. Of course, this process is not the realization, which again can lead to complex situations like financial shortages or the like.
To analyze the models’ behavior systematically and for an inverse problem solution, again, a self-organizing Kohonen map (SOM) is trained. This time, no distinguished regions or clusters can be detected, as was the case with direct interaction when training the map with the simulated time series of $g^{s}$. This is expected when examining Fig. 13 - 16. The differences in the time series are gradual. Still, different phenotypes appear, like exponential convergence, vibrations, or slow convergence, with practically different values for each stakeholder.
@@
  • std(G)
    Convergence detection: The overall standard derivation of all $g^{s}$ is calculated to detect systems convergence and differences between stakeholders $g^{s}$ still present.
  • std(G-gs)
    Bifurcation detection: The ratio between the actual $g$ and the theoretical fixed point Linke2019a $g_{f}=\alpha+\sum\beta_{k}$ is calculated for each stakeholder individually. When the standard deviation of this ratio is calculated, a first impression of whether bifurcations occur can be derived. The SOM is then trained by the mean of this ratio over all stakeholders.
  • Pears2Lin
    Convergence time detection: The absolute value of Pearson Correlation $r$, as calculated from $g^{s}_{t}$ and a straight line ($f(t)=t$), describes whether the $g^{s}_{t}$ is monotonous falling or growing, or if wave-like fluctuations occur, as shows up in Fig. 13.
@@
Figure 11: Trained SOM for mediated interaction between stakeholders using Version II of Varying term III of Eq. 6: a) and b) show the same u-matrix, but the dots, representing the sorted training data are colored following different criteria: In a), the colors are chosen due to the initial conditions. While green dots refer to I1, red dots refer to I2, blue dots refer to I3, and orange dots refer to I4. In b), the colors represent the coupling strength $V$, while darker values refer to weaker coupling.
Fig. 11 has the same u-matrix in plots a) and b). Also, the $g^{s}$ of all cases are best-fit to respective neurons in the same way. The difference between both plots is in the coloring of the dots. While a) colors are according to initial conditions I1 - I4, in b) the colors are chosen according to the coupling strength $V$ of the mediating terms of $\beta^{s,i}$ in Eq. 6.
Differentiating with respect to initial conditions I1 - I4 in Fig. 11 a) shows systematic differences between initial conditions. Only arbitrary $g_{0}$ of I3 and I4 show chaotic transients, with an examples shown in Fig. 15 and Fig. 16). As soon as the $g$ is close to $g_{f}$, the system will behave initially stable. As Eq. 6 changes $\beta$ iteratively, fluctuations are prevented, and the system always shows smooth behavior as already introduced in Fig. 2.
This phenomenon is also strongly influenced by the coupling strength $V$. Fig. 11 b) shows that the SOM also sorts training data according to $V$. The reason for this becomes clear when examining the component planes of the trained SOM, shown in Fig. 12. Columns show the distribution of one feature vector entry and one component over the trained SOM of the three analysis parameters: convergence, bifurcation, and convergence time detectors. By correlating the component planes to the dot distribution, the reason for sorting the $g^{s}$ on the SOM can be found.
In the SOM in Fig. 11 b), the dots in the upper part of the plot are darker, pointing to a weak coupling. The Pears2Lin component plane on the right of Fig. Fig. 12 shows bright values in its upper parts, corresponding to large values of the convergence time detector. Fig. 16 shows that, due to weak coupling, changes of $\beta$ are small, where then only minor changes in $g$ appear. The system is very stable but takes a long time to converge. When the coupling is strong, like in Fig. 13 or 14, the $g^{s}$ approach each other very fast, but might overshoot, leading to vibration behavior. This is detected by the convergence time analysis method Pears2Lin.
@@
Figure 12: Component planes of trained SOM of Fig. 11 of the mediating interaction case, Version II of Varying term III when using Eq. 6. The rows refer to the different time windows. The columns denote the three detection methods of convergence, bifurcation, and convergence time as input feature vectors to the SOM and are displayed as magnitudes of these different features.
Figure 13: Example time Series of $g^{s}$, for each of the s=6 stakeholders, resulting from an IPF model with initial condition I1 and $\alpha^{s}\approx\{0.89,0.87,0.91,0.81,0.92,0.56\}$, $V\approx 0.01$ and $C_{dt}=12$. On the SOM shown in Figure 11, this time series is sorted into the lower left corner.
Figure 14: Example time Series of $g^{s}$, for each of the s=6 stakeholders, resulting from an IPF model with initial condition I2 and $\alpha^{s}\approx\{0.76,0.73,0.80,0.79,0.68,0.84\}$, $V\approx 0.008$ and $C_{dt}=15$. On the SOM shown in Figure 11, this time series is sorted into the lower right corner.
Figure 15: Example time Series of $g^{s}$, for each of the s=6 stakeholders, resulting from an IPF model with initial condition I3 and $\alpha^{s}\approx\{0.89,0.87,0.90,0.81,0.92,0.56\}$, $V\approx 0.003$ and $C_{dt}=2$. On the SOM shown in Figure 11, this time series is sorted into the small island of orange and blue dots, surrounded by red and green dots, on the upper right of the map.
Figure 16: Example time Series of $g^{s}$, for each of the s=6 stakeholders, resulting from an IPF model with initial condition I4 and $\alpha^{s}\approx\{0.86,0.62,0.98,0.91,0.61,0.73\}$, $V\approx 4.7\times 10^{-}5$ and $C_{dt}=2$. On the SOM shown in Figure 11, this time series is sorted into the upper left corner.

### v2

@@ III.1 Single stakeholder without self-adaptation with external disturbance
Five cases are shown in Fig. 3. The top blue curve shows $g^{1}$ over 50 iterations with $I_{t}=0$ without impact. After an initial transient, $g^{1}$ becomes stationary and converges.
@@
The yellow curve shows the reaction to a positive $I_{20}=1$, so an impact of one at a single iteration time 20. The system is disturbed and returns to its converging state. This is expected as such a distortion corresponds to a new initial $g_{0}$ after which convergence is ensured due to the choice of $\alpha=0.56$ for which such a system is stationary.
The green curve shows a stronger impact of $I_{20}=5$. Although the system becomes stationary again, it takes more time to recover.
@@
All these behaviors represent reasonable reactions of a social system to an impact; it might converge to a stationary value, even with a negative impact, or it might break down and leave the stage.
@@ III.2 Single stakeholder with self-adaptation with external impact
To estimate the effect of self-adaptation, the model again with s=1 and a strong impact at iteration time 20 of $I_{20}=10$ is used, but now including self-adaptation, so using varying term II in version I. Fig. 4 compares $g^{s}_{t}$ for both cases, starting at $\alpha=0.56$. In Fig. 3, $\alpha$ = const, while now $\alpha$ is adaptive, therefore time-varying. Also now, $C^{s}=-1$ is used.
The red curve in Fig. 4 is $g^{s}_{t}$ of varying term I version II already shown in Fig. 3. The blue curve is the corresponding curve of varying term II version I, and the adaptive $\alpha$ is shown in gray. Both, $g^{s}_{t}$ and $\alpha_{t}$ are bifurcating but are stable in this respect. The impact at iteration time point 20 disturbs both curves considerably. Still, $g^{s}_{t}$ of version I returns to its bifurcating state much faster than $g^{s}_{t}$ of version II.
@@
Figure 4: Adaptive version I of varying term II compared to non-adaptive version II concerning reaction of external input at iteration time point 20. Red: Version II time series from Fig. 3 with constant $\alpha=0.56$. Blue: Version I time series with adaptive $\alpha$. Compared to Version II, the Version I time series is bifurcating due to adaptivity; still, it is returning to this state faster compared to Version II with static $\alpha$. Gray: Version I adaptive $\alpha$ corresponding to Version I system parameter $g^{s}_{t}$. After the impact on time point 20 $\alpha$ is way above the diverging limit of $\alpha\sim 0.37$, the system remains stable due to adaptation with a maximum $\alpha\sim 0.95$.
@@
All calculations were performed using an initial system value $g^{s}_{0}=0.1$. This ensures $\Im{g_{1}}=0$ for $\alpha_{0}>0.09048$ within the first iterative step, a value of $\alpha$ much lower than expected in the system at all. Using $g_{0}=1$ would lead to $\Im{g_{1}}\neq 0$ for $\alpha_{0}\geq e$. Therefore, any adaptation algorithm would no longer be able to converge $\alpha_{0}<0.3679$, and extended stability by including any adaptation algorithm could not be tested. Using any $g_{0}$ is artificial in any case, as the initial time point of a social or political situation is arbitrarily set by an investigator while the stakeholders taking part in the social action most often existed before the chosen model starting time point.
@@
Fig 5 shows the three cases for I = 0, so without adaptation, I = 1, and I = 10. The adaptation constant was varied from -3 $\leq C^{s}\leq 7$ and the initial $\alpha_{0}$ like 0.1961 $\leq\alpha_{0}\leq$ 10.
In the top plot of Fig. 5, colors denote converging cases of $C^{s}$ vs. $\alpha_{0}$ combinations where the color corresponds to the mean converged value. A tail appears starting from $\alpha_{0}$ = 10 and 0 $\leq C^{s}\sim$2 which is getting narrower towards $\alpha_{0}\sim 0.7$ and $C^{s}\sim$ 2. Therefore, choosing a suitable $C^{s}$ allows for much higher $\alpha_{0}$ to converge compared to the non-adaptive Version II with a minimum of $\alpha\sim 0.37$. Indeed, with decreasing $\alpha_{0}$, the range of $C^{s}$ for convergence is getting smaller, as expected. It is reliable that the convergence value is smaller towards higher $\alpha$ displayed by the lighter color as the IPF converges against $\alpha^{s}+\sum{\beta^{s,i}}$, which has already been shown analytically for the nonadaptive caseLinke2019a .
Nevertheless, a cohesive area in which the IPF converges for all enclosed parameter values can always be found and is never fragmented in the sense of non-converging islands. Thus, convergence is ensured within the displayed regions.
@@
An unsystematic convergence area also exists for $C^{s}<-2$, which is quite speckled. Stakeholders in this region must be modeled very carefully. As the stable regions are narrow, small changes in the initial condition may drastically change the results.
@@
III.3 Multiple stakeholders with fixed impact and no external disturbance
Adding the interaction between multiple stakeholders leads to a temporal description of the system parameter $g^{s}$ for all stakeholders in the process. This is implemented by the interaction parameter $\beta^{s,i}_{t}$ of stakeholders i impacting on stakeholder s. First, fixed impact is discussed, which means including the Version I of Varying term III as shown in Tab. 1.
@@ III.3 Multiple stakeholders with fixed impact and no external disturbance
In this simulation, we assume all stakeholders are already present in the system before the simulation starts at $t=0$. Still, couplings $\beta^{s,i}$ are often not known or hard to determine, which leads to the simplification of uncoupled stakeholders before $t=0$. Thus Eq. 4 and 8 are valid $\forall\ t>=0$. $\forall\ t<0$ the system can be described by the IPF with $\beta^{s,i}=0\forall\ s,i$. As we assume the system to be running for a sufficiently long time, all stakeholders have already converged to their fixed point. When assuming that no $\beta$ are involved before $t=0$ the system is at the fixed point $g^{s}_{f}=\alpha^{s}$ according to Eq. 4 and thus $g^{s}_{t}=\alpha^{s}\ \forall t<0$. Then, all necessary initial values of the simulation are set. Further, the time delay $d^{s,i}=i$ directly depends on the couped stakeholder.
In the following, IPFs with $N=6$ stakeholders are used throughout, as taken from the Frappant example. Each IPF is iterated for 500 steps using Equations 4 and 8. $d^{s,i}=1$ is used throughout. By randomly varying $\alpha^{s}$, IPFs were calculated until 1000 non-diverging time series were obtained. This amount was reached after 3523 attempts. Thereby, $\alpha^{s}$ was chosen within a range of $\alpha^{s}\in[0.2,1]$. Also, the coupling constant $W^{s}_{i}=W$ was fixed for all stakeholders but varied for each IPF again randomly in the range of $W\in[0.01,1/6]$.
@@
Figure 6: Trained SOM of 1000 IPF cases of six stakeholders using fixed interaction with Version I of Varying terms II. The background color is the u-matrix displaying similarities (dark colors) or differences (light colors) of neighbouring neurons. Therefore, the yellow ridges are boundaries between clusters present in darker regions. Each dot is an IPF $g^{s}$ iteration’, which is placed according to its best fit to the feature vectors of the trained SOM. The colors of the dots refer to the coupling strength $W\in[0.01,1/6]$; lighter/darker colors correspond to weaker/stronger coupling. For each of the clusters or regions I to VIII, an assigned time series $g^{s}_{t}$ is picked and plotted aside the SOM as a typical example of stakeholders $g^{s}$ development of the last 20 values of the 500 iteration steps for the respective regions I to VIII.
Fig. 6 shows the trained SOM for Version I of Varying term III. The background color is the so-called u-matrix, displaying similarities and differences of neighboring neurons. Therefore, the Euclidean distances of a neuron to its neighbors are calculated in the feature space. Then, the mean value of these distances allows an evaluation of how far away the neurons are located in the feature space and thus, how different they are.
In Fig. 6, lighter yellow colors have strong differences, and darker blue colors have small differences. Therefore, the yellow ridges are boundaries between blue regions of very similar neurons. The dots represent the 1000 IPF cases the map was trained with, and they are positioned after training at the best-fit neurons for each IPF. The colors of the dots show varying coupling strengths $W$, where lighter colors are weaker and darker colors have stronger coupling.
@@
Figure 7: Example IPF time series of region I of Fig. 6 of $g^{s}$ converging for all N=6 stakeholders, calculated from Eq. 8. The initial adaptation vector was $\alpha^{s}\approx\{0.98,0.71,0.79,0.75,0.55,0.77\}$ and $W\approx 0.011$. Therefore, the largest region I of all eight regions is that of convergence of all stakeholders.
@@
A more systematic view of regions II - VII can be obtained from Fig. 10. Here, for each node of the SOM, the last ten values of the time series $g^{s}_{t}$ are inspected in terms of the number of different values. Values are assumed to be different when they differ more than 5 % from each other. One can clearly distinguish the stable region I and the chaotic region VIII. Furthermore, II - VII, while looking rather similar in Figure 6, are found to differ in terms of the order of bifurcation.
Figure 10: Number of different values $g^{s}_{t}$ when inspecting the last ten values of each stakeholder time series. Shown for every neuron of the trained SOM of Fig. 6. Values are assumed to be different when they differ more than 5 % from each other. Dark/light regions refer to low/high numbers of different values. Region I clearly has the lowest values. As it is converging throughout, no fluctuations are detected. Also, the chaotic region VIII has the highest values. Still, region II - VII can now be differentiated with respect to the number of different values and, therefore, the order of bifurcation.
Using only version I of varying terms III of fixed interaction between stakeholders shows all kinds of possible scenarios for city planning. These include
  • Stationary, Convergence: All stakeholders arrive at their desired value.
@@
So, the inverse problem of finding parameters and therefore, political actions to arrive at a desired output for all stakeholders is solved using the proposed method. Still, a second interaction process will appear between stakeholders, that of adaptive interactions discussed below.
III.4 Multiple stakeholders with adaptive impact and no external disturbance
Along the method of fixed interaction discussed above, for adaptive interaction again IPFs for $N=6$ stakeholders with 500 iteration steps each were calculated, using Eq. 4 and now Version II of Varying terms III of Eq. 9.
@@ III.4 Multiple stakeholders with adaptive impact and no external disturbance
  • I1: Same as the previous modeling example of fixed impact, it is assumed that no coupling is present $\forall t<0$ and that the system has already settled itself, and thus $g^{s}_{t}=\alpha^{s}\ \forall t<0$. At $t=0$, suddenly, all stakeholders are coupled to each other.
  • I2: Similar to I1, but as the system is supposed to be uncoupled $\forall t<0$, it is not realistic that all stakeholders are suddenly coupled to each other. The IPF models the impact of the coupled stakeholders by the damped impact of preceding system states $g^{s}_{t-d^{s,i}}$. As coupling is only present for $t\geq 0$, previous system states can not act on future ones. Thus, the different $\beta^{s,i}_{t}$ are only taken into account if $t\geq d^{s,i}$.
  • I3: It is assumed that the system is not settled at the start. Thus $g^{s}_{t}$ may have an arbitrary value $\forall t<0$. Here, for simplicity, we chose $g^{s}=1$. Like in I1 it is assumed that all stakeholders are suddenly coupled to each other at $t=0$.
@@
The resulting time series are shown in Fig. 13 to 16 as examples for each of the four initial conditions. They are very much different from those of the fixed interactions discussed above. Bifurcations and chaos are rarely present, rather damped oscillations appear. Also, due to the mutual adaptation of $\beta^{s,i}_{t}$, all coupled IPFs converge to the same fixed point. This fixed point might be reached in an exponential development, like shown in Fig.15, it might show vibrations before convergence, like in Fig. 13 - 14, or it might show very slow convergence, so basically each stakeholder keeps its own state.
An example is a legally binding land-use plan process (B-Plan Verfahren) in which all stakeholders negotiate over time on a planning of a area. The result is planning certainty. Due to negotiations, an adaptive and self-reference process is leading to perfect convergence of all stakeholders needs. Of course, this process is not the realization, which again can lead to complex situations like financial shortages or the like.
To analyze the models’ behavior systematically and for an inverse problem solution, again, a self-organizing Kohonen map (SOM) is trained. This time, no distinguished regions or clusters can be detected, as was the case with fixed interaction when training the map with the simulated time series of $g^{s}$. This is expected when examining Fig. 13 - 16. The differences in the time series are gradual. Still, different phenotypes appear, like exponential convergence, vibrations, or slow convergence, with practically different values for each stakeholder.
@@
  • Convergence detection:
    The overall standard derivation of the system states of all stakeholders in a certain time window $g^{s}_{t}\forall s\in[0,6],t\in[t_{n},t_{m}]$ is calculated to detect systems convergence and if differences between stakeholders $g^{s}_{t_{n}}$ is still present.
  • Oscillation detection:
    The ratio between the actual $g$ and the theoretical fixed point Linke2019a $g_{f}=\alpha+\sum\beta_{k}$ is calculated for each stakeholder individually. When the standard deviation of this ratio is calculated, a first impression of whether damped oscillations occur can be derived. The SOM is then trained by the mean of this ratio over all stakeholders.
  • Transient evaluation:
    The absolute value of Pearson Correlation $r=\frac{conv(X,Y)}{\sigma_{X}\sigma_{Y}}$ is calculated. With $conv(X,Y)$ being the covariance of the two Vectros $X,Y$ and $\sigma_{X},\sigma_{Y}$ being the standard derivation of the two Vectors.
    If $X$ is the time series of system states of a stakeholder $g^{s}_{t}$ and $Y$ a straight line $f(t)=t$, $r^{s}$ describes whether the $g^{s}_{t}$ is monotonous falling or growing, or if wave-like fluctuations occur, as shows up in Fig. 13. By calculating the mean value of the absolute value of $r^{s}$ over all stakeholders the transient behavior can be evaluated.
@@
Figure 11: Trained SOM for adaptive interaction between stakeholders using Version II of Varying term III of Eq. 9: a) and b) show the same u-matrix, but the dots, representing the sorted training data are colored following different criteria: In a), the colors are chosen due to the initial conditions. While green dots refer to I1, red dots refer to I2, blue dots refer to I3, and orange dots refer to I4. In b), the colors represent the coupling strength $V$, while darker values refer to weaker coupling.
Fig. 11 has the same u-matrix in plots a) and b). Also, the $g^{s}$ of all cases are best-fit to respective neurons in the same way. The difference between both plots is in the coloring of the dots. While a) colors are according to initial conditions I1 - I4, in b) the colors are chosen according to the coupling strength $V$ of the adaptive terms of $\beta^{s,i}$ in Eq. 9.
Differentiating with respect to initial conditions I1 - I4 in Fig. 11 a) shows systematic differences between initial conditions. Only arbitrary $g_{0}$ of I3 and I4 show chaotic transients, with an examples shown in Fig. 15 and Fig. 16). As soon as the $g$ is close to $g_{f}$, the system will behave initially stationary. As Eq. 9 changes $\beta$ iteratively, fluctuations are prevented, and the system always shows smooth behavior as already shown in Fig. 2.
This phenomenon is also strongly influenced by the coupling strength $V$. Fig. 11 b) shows that the SOM also sorts training data according to $V$. The reason for this becomes clear when examining the component planes of the trained SOM, shown in Fig. 12. Columns show the distribution of one feature vector entry and one component over the trained SOM of the three analysis parameters: convergence and oscillation detectors as well as the transient evaluation. By correlating the component planes to the dot distribution, the reason for sorting the $g^{s}$ on the SOM can be found.
In the SOM in Fig. 11 b), the dots in the upper part of the plot are darker, pointing to a weak coupling. The transient evaluation component planes on the right of Fig. 12 show bright values in their upper parts, corresponding to large values of the transient evaluation and thus linear behavior. Fig. 16 shows that, due to weak coupling, changes of $\beta$ are small, such that only minor changes in $g$ appear. The system is very stationary but takes a long time to converge. When the coupling is strong, like in Fig. 13 or 14, the $g^{s}$ approach each other very fast, but might overshoot, leading to vibration behavior. This is detected by the transient evaluation method.
@@
Figure 12: Component planes of trained SOM of Fig. 11 of the adaptive interaction case, Version II of Varying term III when using Eq. 9. The rows refer to the different time windows. The columns denote the three detection methods of convergence, oscillation, and transient evaluation as input feature vectors to the SOM and are displayed as magnitudes of these different features.
Figure 13: Example time series of $g^{s}$, for each of the N=6 stakeholders, resulting from an IPF model with initial condition I1 and $\alpha^{s}\approx\{0.89,0.87,0.91,0.81,0.92,0.56\}$, $V\approx 0.01$ and $C_{dt}=12$. On the SOM shown in Figure 11, this time series is sorted into the lower left corner.
Figure 14: Example time series of $g^{s}$, for each of the N=6 stakeholders, resulting from an IPF model with initial condition I2 and $\alpha^{s}\approx\{0.76,0.73,0.80,0.79,0.68,0.84\}$, $V\approx 0.008$ and $C_{dt}=15$. On the SOM shown in Figure 11, this time series is sorted into the lower right corner.
Figure 15: Example time series of $g^{s}$, for each of the N=6 stakeholders, resulting from an IPF model with initial condition I3 and $\alpha^{s}\approx\{0.89,0.87,0.90,0.81,0.92,0.56\}$, $V\approx 0.003$ and $C_{dt}=2$. On the SOM shown in Figure 11, this time series is sorted into the small island of orange and blue dots, surrounded by red and green dots, in the upper right of the map.
Figure 16: Example time series of $g^{s}$, for each of the N=6 stakeholders, resulting from an IPF model with initial condition I4 and $\alpha^{s}\approx\{0.86,0.62,0.98,0.91,0.61,0.73\}$, $V\approx 4.7\times 10^{-}5$ and $C_{dt}=2$. On the SOM shown in Figure 11, this time series is sorted into the upper left corner.